Question 6 !! Thanks
D2L Bright space 5. Order the following functions according to their order of growth (from the lowest to n!, nlgn, 81g (n 10) 10 23, 32n, n 101gn 6. Prove that a lg(n c) O(lg n), for every fixed k 0, a 0 and c >0. 7. Determine the complexities of the following recursive functions, where c 0 is the 2 (a) Consider the recursive function, use the Big O notation to determine its com T(n) (n/2) c n, if n 1;
Answer #1

Ans 6 : g(n) = lg(n)

f(n)= a + lg(n^k + c)

a+lg(nk *(1+c/nk))

a+lg(nk)+lg(1+c/nk)

a+ klg(n)+lg(1+c/nk)

as a,k,e all are >0 , hence we can neglect the value lg(1+c/nk) as very small

hence the upper bound for the above function in O(lg n) (1)

similarly f(n) grows as fast as g(n), thtrefore , f(n)= Ω(g(n)) (2)

from above two points we can conclude that f(n)=Θ(g(n))